Circular No. 7134

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999be AND 1999bf P. Nugent, Lawrence Berkeley National Laboratory, reports the discov-

ery of two supernovae on unfiltered CCD images on behalf of the Supernova Cosmology Project (cf. IAUC 7125, 7128) and the Spacewatch program at the University of Arizona (which includes R. S. McMillan, T. Gehrels, J. A. Larsen, J. L. Montani, J. V. Scotti, N. Danzl, and A. Gleason):

SN 1999 UT α_{2000} δ_{2000} Mag. Offset
1999be Mar. 10 8^{h}23^{m}38.^{s}60 +18^{o}18′20′′.4 15.8 12′′ W, 24′′ N
1999bf Mar. 10 11 49 52.28 − 0 41 19.9 20.8  ?

A CCD spectrum by G. Aldering at the Cerro Tololo 4-m telescope on Mar. 25 UT showed SN 1999be to be a type-Ia supernova, ∼ 1 week after peak brightness, at a recessional velocity of 5650 km/s, similar to that of the host galaxy, CGCG 089-013. SN 1999bf has no apparent host on the discovery images (limiting mag ∼ 22); it was also observed by Aldering the same night and was found to be a type-Ia supernova near maximum light at a redshift of 0.24. Finding charts for the supernovae can be obtained

from http://panisse.lbl.gov/nearsearch/fcharts/index.html.

SUPERNOVA 1999bd IN ANONYMOUS GALAXY B. Skiff, Lowell Observatory, reports that the Digitized Sky Survey

shows a very faint galaxy of nominal red magnitude 19.0 whose USNO-A2.0 catalogue position (α = 9^{h}30^{m}29^{s}.10, δ = +16^{o}26′06′′.8, equinox 2000.0) is 1′′.4 from the supernova position reported on IAUC 7133.

COMET P/1999 D1 (HERMANN) Ephemeris extension to IAUC 7112 from orbital elements on MPC

33858 (P = 13.4 yr):

1999 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 23 11^{h}43.^{m}07 −10^{o}23.^{′}1 0.704 1.692 168.^{o}7 6.^{o}6 16.5
Apr. 2 11 32.53 −13 54.8 0.746 1.720 160.2 11.3 16.7
12 11 25.23 −16 42.6 0.810 1.754 150.8 16.2 17.0
22 11 22.05 −18 52.3 0.892 1.794 142.1 20.1 17.3
May 2 11 23.06 −20 34.4 0.989 1.839 134.2 23.1 17.6
12 11 27.79 −21 58.2 1.100 1.889 127.1 25.3 18.0
22 11 35.67 −23 11.2 1.222 1.943 120.6 26.6 18.3
1999 March 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green